# Causal Lifting of Neural Representations: Zero-Shot Generalization for Causal Inferences

### v2

@@ 3 Causal Lifting of Neural Representations
In the previous Section we presented and discussed the PPCI problem and its peculiar challenges to learn a valid outcome model. In this Section, we introduce a new representation learning constraint to prevent such generalization challenges, and corresponding implementation, tailored for our application of interest.
@@
Given a PPCI problem $\mathscr{P}$ with standard causal identification assumptions, and a predictor $g=h\circ\phi:\mathcal{X}\rightarrow\mathcal{Y}$. The encoder $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$ satisfies the Causal Lifting constraint when:
@@
where $Z=[T,\tilde{W}]$, and $\tilde{W}\subseteq W$ is a valid adjustment set.
The Causal Lifting constraint is a condition on the model encoder to prevent the possibility to leverage unnecessary spurious outcome information from the experimental settings.
@@
Given two similar PPCI problems $\mathscr{P}$ and $\mathscr{P}^{\prime}$ with standard causal identification assumptions. Let $g=h\circ\phi:\mathcal{X}\rightarrow\mathcal{Y}$ be an outcome model for $\mathscr{P}^{\prime}$ with $h$ Bayes-optimal, and assume that the representation transfers, i.e., $\mathbb{P}(\phi(X)|Y)=\mathbb{P}^{\prime}(\phi(X)|Y)$. Then the Causal Lifting constraint on $\mathscr{P}$ implies $g$ is valid on $\mathscr{P}$.
Theorem 3.1 offers a sufficient condition to transfer model validity from a training experiment to a target experiment by assuming the Causal Lifting constraint holding on target. It combines the strengths of modern deep learning models, able to extract sufficient representations on training tasks, i.e., $I(Y|\phi(X))=I(Y|X)$, together with efficient fitting on limited data, i.e., Bayes Optimal classifier assumption. We remark that the assumption that the representation transfers is reasonable, since we have assumed that an invariant annotation protocol exists. In practice, the Causal Lifting constraint is required to hold on the target experiment, where we cannot get any guarantees without labels. However, referring to the generalization challenges examples in Section 2:
@@
Despite the unverifiability of this assumption, we practically rely on the hope that transferring such property is an easier generalization task than transferring conditional calibration. We then propose to enforce the causal lifting constraint during training/fine-tuning on the treatment and the valid adjustment set for the target experiment, when also observed. If not observed but constant in the training experiment, the condition is satisfied by default (no signal for the model to learn such experimental settings). As pre-trained models trained on huge generic data are very capable of detecting all visible information that can correlate with a prediction target, the causal-lifting constraint has to be enforced explicitly during transfer learning, forcing the artificial annotation process to be “blind”. Similar to the Conditional Calibration assumption, if the outcome is predictable well from the measurement, enforcing the causal lifting constraint on more experimental settings than necessary does not hurt.
@@ 3.1 Deconfounded Empirical Risk Minimization
The Causal Lifting constraint is a conditional independence condition, and as such, enforcing it during training is a well-known problem in the Representation Learning literature. Different direct and indirect approaches have already been proposed, and in Section 4 we overview the major paradigms. Here, we propose a simple implementation that is tailored to common assumptions in scientific experiments (and that are true in ISTAnt), i.e., that experimental settings are low-dimensional and discrete (Pearl et al., 2000; Rosenbaum et al., 2010). Other more general implementations are certainly possible if such assumptions do not hold, but this is beyond the scope of this paper. In detail, we leverage a resampling approach (Kirichenko et al., 2022; Li and Vasconcelos, 2019), reweighting the training distribution to simulate a distribution that has no correlation between outcome and experimental settings, while keeping the annotation mechanism invariant. We define the manipulated distribution of $Z^{\star}$,$Y^{\star}$ as function of $Z,Y$ and joint distribution:
@@
Finally, we remark that when Condition 9 doesn’t hold, it is not possible to specify a ‘disentangled’ joint distribution without ignoring the experimental settings, i.e., $\mathbb{P}(Z^{\star}=z)=0$, where the conditional outcome’s support is strictly contained in the marginal outcome’s support on the training distribution. Our manipulated distribution will still consider these samples, but reducing their weight with respect to the predictivity of the observed experimental setting (approximately $\propto\text{Var}(Y|Z=z)$). Therefore, some spurious correlations may still be learned, but this is now in a trade-off with ignoring a potentially substantial part of the reference sample. While this does not apply to our application of interest, tailored modifications of our joint distribution can be proposed case-by-case.
@@ 4 Related Works
Prediction-Powered Causal Inference
The factual outcome estimation problem for causal inference from high-dimensional observations was first introduced by Cadei et al. (2024). We extended the problem to generalization from similar experiments, motivated by practical desiderata, e.g., in experimental ecology. In our paper, we formalize what they describe as “encoder bias”, and our DERM is the first proposal to their call for “new methodologies to mitigate this bias during adaptation”, although we consider an additional OOD component. Demirel et al. (2024); Wang et al. (2024) already attempted to discuss some generalization challenges in a PPCI problem, but with unrealistic assumptions. They ignored any high-dimensional measurement of the outcome of interest and assumed that the experimental settings alone are sufficient for factual outcome estimation together with support overlapping, i.e., without generalization, making the model too application-specific and ignoring any connection with representation learning. Let’s further observe that their framework is a special case of ours when $X$ is low-dimensional, and the target experiment is in-distribution. In contrast to the classic Prediction-Powered Inference (PPI) (Angelopoulos et al., 2023a, b), which improves estimation efficiency by imputing unlabeled in-distribution data via a predictive model, and recent causal inference extensions (Wang et al., 2024; De Bartolomeis et al., 2025; Poulet et al., 2025) relying on counterfactual predictions, PPCI focuses on imputing missing factual outcomes to generalize across similar experiments.
@@
Learning representations invariant to certain attributes is a widely studied problem in machine learning (Moyer et al., 2018; Arjovsky et al., 2019; Gulrajani and Lopez-Paz, 2021; Krueger et al., 2021; Yao et al., 2024a). In some sense, we wish to learn representations of $X$ that can predict the outcome $Y$, but are invariant to the experimental variables $Z$. In agreement with Yao et al. (2024a), we combine a sufficiency objective with our causal lifting constraint in the representation space, enforcing a suitable representation invariance. Several alternative approaches can be considered to enforce such conditional independence constraints: (i) Conditional Mutual Information Minimization (Song et al., 2019; Cheng et al., 2020; Gupta et al., 2021), (ii) Adversarial Independence Regularization such as Louizos et al. (2015) which modifies variational autoencoder (VAE) architecture in Kingma (2014) to learn fair representations that are invariant sensitive attributes, by training against an adversary that tries to predict those variables, (iii) Conditional Contrastive Learning such as Ma et al. (2021) whereby one learns representations invariant to certain attributes by optimizing a conditional contrastive loss, (iv) Variational Information Bottleneck methods where one learns useful and sufficient representations invariant to a specific domain (Alemi et al., 2016; Li et al., 2022). We chose our DERM procedure as it matches well our application of interest, but other applications may prefer different conditional independence implementations.
@@
In the broader context of causal representation learning methods (Schölkopf et al., 2021), our proposal largely focuses on representation learning applications to causal inference: learning representations of data that make it possible to estimate causal estimands. We find it in contrast with most recent works in causal representation learning, which uniquely focused on complete identifiability of all the variables or blocks, see Varici et al. (2024); von Kügelgen (2024) for recent overviews targeting general settings. The main exceptions are Yao et al. (2024a, b). The former leverages domain generalization regularizers to debias treatment effect estimation in ISTAnt from selection bias. However, their proposal is not sufficient to prevent confounding when no data from the target experiment is given. The latter uses multi-view causal representation learning models to model confounding for adjustment in an observational climate application. In our paper, we also discuss conditions for identification, but we focus on a specific causal estimand, as opposed to block-identifiability of causal variables. As opposed to virtually all existing work in causal representation learning, our perspective offers clear evaluation and benchmarking potential on real world scientific experiments – even in theoretically under-specified settings via the accuracy of the causal estimate.
@@ 5 Experiments
In pursuit of scientific validity, we evaluate our method (DERM) on the ISTAnt dataset (Cadei et al., 2024), and additional synthetic experiments controlling for the true causal effect.
Zero-Shot Generalization on ISTAnt.   We performed a similar experiment to the ISTAnt experiment on our recording platform, considering three different treatments and collecting 44 annotated 30-minute videos, and we leveraged such annotated videos to fine-tune a pretrained model generalizing zero-shot prediction powered ATE inference on ISTAnt by AIPW. We considered the AIPW estimation on the (true) ISTAnt human annotated outcomes as the ground truth ATE (approximately $+40$s in grooming time per video). Further details on the modeling choices, hyper-parameter, and fine-tuning are discussed in Appendix B. Figure 3 shows the differences in filming setup among the two experiments.
Figure 2: 0-shot PP-ATE inference on ISTAnt dataset from our similar experiment, varying method and pre-trained encoder. Reported the 95% confidence intervals estimated via AIPW asymptotic normality and compared against the same inference on the annotated outcomes. In green the correct confidence intervals (overlapping 33+%) in red the incorrect.
@@
CausalMNIST.   We replicate the analysis on a colored manipulations of the MNIST dataset, e.g., estimating the effect of the background color or pen color on the digit value, allowing complete control of the causal effects and uncertainty quantification by Monte Carlo simulations. We train a convolutional neural network on an experiment and we test zero-shot PPCI on four similar observational studies, varying the treatment effect mechanism and experimental settings distribution, challenging the full support assumption, i.e., Equation 9. A full description of the data-generating processes and analysis is reported in Appendix C.
The results are reported in Table 1, clearly showing the DERM potential to solve unsolvable instances for both ERM and invariant training (see soft shifts columns), while suffering when challenging its assumptions (see hard shifts columns).
Table 1: Zero-shot PPCI results on CausalMNIST varying the target experiment population (in-, out-of-distribution), the treatment effect mechanism (linear, non-linear) and challenging the full-support guarantees (soft, hard shifts). We report the ATE estimation bias and standard deviation over 50 repetitions.
@@ 6 Conclusion
In this paper, we introduced the challenge of solving PPCI problems zero-shot. We are motivated by scientific applications, where predictors can be trained on existing data collected in prior experiments and power valid causal conclusions in new experiments without further training. Theoretically, we developed the first identification and estimation results for PPCI, alongside a causal lifting constraint that transfers causal validity under standard yet idealized assumptions. In practice, we have offered an implementation of the causal lifting constraint with our Deconfounded Empirical Risk Minimization, and showed that it achieves zero-shot identification of the ATE on the ISTAnt dataset (Cadei et al., 2024) by leveraging (our) similar ecological experiment. Overall, our work offers a paradigm shift from the traditional causal representation learning literature (Schölkopf et al., 2021), paving the way toward learning representations that power downstream causal estimates on real-world data. We think that this has the potential to play a critical role for deep learning models to accelerate the analysis of scientific data and ultimately new discoveries.
The main limitation of this work is that, in practice, it is not possible to verify whether either the conditional calibration property or the causal lifting constraint actually holds on the target population without annotations. Beyond our key assumption that the representation transfers, we also did not discuss the model convergence on finite training datasets, so even if enforcing the causal lifting on training would transfer, we may still suffer from finite sample effects. Another limitation is neglecting several settings that are, potentially, practically relevant, e.g., not overlapping effect modifiers among similar experiments, or the violation of Condition 9. As our target experiments did not have these problems, we left them for future work. Finally, we rest on the assumption that the underlying causal inference problem is statistically identifiable, but this is necessary and can be ensured with a well-designed experimental protocol. Despite these limitations, we hope our work can inspire a new generation of practical causal representation learning methodologies and create opportunities for more systematic and practically relevant benchmarking.
@@ Acknowledgments
We thank the Causal Learning and Artificial Intelligence group at ISTA for the continuous feedback on the project and valuable discussions. We thank the Social Immunity group at ISTA, particularly Jinook Oh, for the annotation program and Michaela Hoenigsberger for supporting our ecological experiment. Riccardo Cadei is supported by a Google Research Scholar Award and a Google Initiated Gift to Francesco Locatello. This research was funded in part by the Austrian Science Fund (FWF) 10.55776/COE12). It was further in part supported by the ISTA Interdisciplinary Project Committee for the collaborative project “ALED” between Francesco Locatello and Sylvia Cremer. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.

### v3

@@ 3 Causal Lifting of Neural Representations
In the previous Section, we introduced and discussed PPCI and its peculiar challenges to learn a valid outcome model from training data out-of-distribution. In this Section, we introduce a new representation learning constraint to prevent such generalization challenges, and a corresponding implementation, tailored for our application of interest.
@@
Given a PPCI problem $\mathscr{P}$, with identifiable ATE (given the ground-truth annotations), and a predictor $g=h\circ\phi:\mathcal{X}\rightarrow\mathcal{Y}$. The encoder $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d}$ satisfies the Causal Lifting constraint if:
@@
where $Z=[T,\tilde{W}$], and $\tilde{W}\subseteq W$ is a valid adjustment set for the ATE identification.
The Causal Lifting constraint is a condition on the model encoder to prevent the possibility of leveraging unnecessary spurious outcome information from the visible experimental settings.
@@
Given two similar PPCI problems $\mathscr{P}$ and $\mathscr{P}^{\prime}$, with identifiable ATE. Let $g=h\circ\phi:\mathcal{X}\rightarrow\mathcal{Y}$ be an outcome model for $\mathscr{P}^{\prime}$ with $h$ Bayes-optimal, and assume that the representation transfers, i.e., $\mathbb{P}(\phi(X)|Y=y)=\mathbb{P}(\phi(X^{\prime})|Y^{\prime}=y)$. Then the Causal Lifting constraint on $\mathscr{P}$ implies $g$ is valid on $\mathscr{P}$.
Theorem 3.1 offers a sufficient condition to transfer model validity from a training experiment to a target experiment by assuming the Causal Lifting constraint holds on target; see proof in Appendix A. It combines the strengths of modern deep learning models, able to extract sufficient representations on training tasks, i.e., $I(Y|\phi(X))=I(Y|X)$, together with efficient fitting on limited data, i.e., Bayes Optimal classifier assumption. We remark that the assumption that the representation transfers is reasonable, since invariant annotation protocol commonly exists. In practice, the Causal Lifting constraint is required to hold on the target experiment, where we cannot get any guarantees without labels. However, referring to the generalization challenges examples in Section 2:
@@
Despite the unverifiability of this assumption, we practically rely on the bet that transferring such property is an easier generalization task than transferring conditional calibration directly. We then propose to enforce the causal lifting constraint during training/fine-tuning on the treatment and the valid adjustment set for the target experiment, when also observed. If not observed but constant in the training experiment, the condition is satisfied by default (no signal for the model to learn such experimental settings). As pre-trained models over huge generic corpora get capable of detecting all visible information that can correlate with a prediction target, the causal-lifting constraint has to be enforced explicitly during transfer learning, forcing the artificial annotation process to be “blind” with respect to certain information. Similar to the Conditional Calibration assumption, if the outcome is well predictable from the measurement, enforcing the causal lifting constraint on more experimental settings than necessary does not hurt.
@@ 3.1 Deconfounded Empirical Risk Minimization
The Causal Lifting constraint is a conditional independence condition, and as such, enforcing it during training is a well-known problem in the Representation Learning literature. Different direct and indirect approaches have already been proposed, and in Section 4 we overview the major paradigms. Here, we propose a simple implementation that is tailored to common assumptions in scientific experiments (and that are true in ISTAnt), i.e., that experimental settings are low-dimensional and discrete (Pearl et al., 2000; Rosenbaum et al., 2010). Other more general implementations are certainly possible if such assumptions do not hold, but this is beyond the scope of this paper. We leverage a resampling approach (Kirichenko et al., 2022; Li and Vasconcelos, 2019), reweighting the training distribution to simulate a distribution that has no correlation between outcome and experimental settings, while keeping the annotation mechanism invariant. We define the manipulated distribution of $Z^{\star}$$Y^{\star}$ as function of $Z,Y$ and their joint distribution:
@@
Finally, we remark that when the condition in Equation 8 doesn’t hold, it is not possible to specify a ‘disentangled’ joint distribution without ignoring the experimental settings, i.e., $\mathbb{P}(Z^{\star}=z)=0$, where the conditional outcome’s support is strictly contained in the marginal outcome’s support on the training distribution. Our manipulated distribution will still consider these samples, but reducing their weight with respect to the predictivity of the observed experimental setting (approximately $\propto\text{Var}(Y|Z=z)$). Therefore, some spurious correlations may still be learned, but this is now in a trade-off with ignoring a potentially substantial part of the reference sample. While this does not apply to our application of interest, tailored modifications of our joint distribution can be proposed case-by-case.
@@ 4 Related Works
Prediction-Powered Inference
The factual outcome estimation problem for causal inferences from complex observations was first introduced by Cadei et al. (2024). We provide here its first formulation and corresponding foundational discussions and a valid approach to the problem, i.e., causal lifting pretrained neural representation. Particularly, we formalize what they describe as “encoder bias”, and our DERM is the first proposal to their call for “new methodologies to mitigate this bias during adaptation” under similar yet different fine-tuning distribution. Demirel et al. (2024); Wang et al. (2024) already attempted to discuss similar generalization challenges with artificial predictions for causal inferences. However, they both ignored the main problem motivation to retrieve the latent outcomes from complex measurements, and not just weak signal from limited experiment settings, relying instead on unrealistic distributional assumptions. They ignored any entangled and high-dimensional measurement of the outcome of interest and assumed that a few experimental settings alone are sufficient for factual outcome estimation, together with support overlapping, i.e., without generalization, making any model application-specific and ignoring any connection with general-purpose foundational models. Let’s further observe that their framework is a special case of ours when $X$ is interpretable and low-dimensional, and the target experiment is in-distribution. In contrast to the classic Prediction-Powered Inference (PPI) (Angelopoulos et al., 2023a, b), which improves estimation efficiency by imputing unlabeled in-distribution data via a predictive model, and recent causal inference extensions (Wang et al., 2024; De Bartolomeis et al., 2025; Poulet et al., 2025) relying on counterfactual predictions, PPCI focuses on imputing missing factual outcomes to generalize across similar experiments, yet out-of-distribution.
@@
Learning representations invariant to certain attributes is a widely studied problem in machine learning (Moyer et al., 2018; Arjovsky et al., 2019; Gulrajani and Lopez-Paz, 2021; Krueger et al., 2021; Yao et al., 2024a). Practically, we aim to learn representations of $X$ that can predict the outcome $Y$, but are invariant to the experimental variables $Z$. In agreement with Yao et al. (2024a), we combine a sufficiency objective with our causal lifting constraint in the representation space, enforcing the suitable representation invariance. Several alternative approaches can be considered to enforce such conditional independence constraints: (i) Conditional Mutual Information Minimization (Song et al., 2019; Cheng et al., 2020; Gupta et al., 2021), (ii) Adversarial Independence Regularization such as Louizos et al. (2015) which modifies variational autoencoder (VAE) architecture in Kingma (2014) to learn fair representations that are invariant sensitive attributes, by training against an adversary that tries to predict those variables, (iii) Conditional Contrastive Learning such as Ma et al. (2021) whereby one learns representations invariant to certain attributes by optimizing a conditional contrastive loss, (iv) Variational Information Bottleneck methods where one learns useful and sufficient representations invariant to a specific domain (Alemi et al., 2016; Li et al., 2022). We chose our DERM procedure as it matches well our application of interest, but other applications may prefer different conditional independence implementations.
@@
In the broader context of causal representation learning methods (Schölkopf et al., 2021), our proposal largely focuses on representation learning applications to causal inference: learning representations of data that make it possible to estimate causal estimands. We find it in contrast with most recent works in causal representation learning, which uniquely focused on complete identifiability of all the variables or blocks, see Varici et al. (2024); von Kügelgen (2024) for recent overviews targeting general settings. The main exceptions are Yao et al. (2024a, b). The former leverages domain generalization regularizers to debias treatment effect estimation in ISTAnt from selection bias. However, their proposal is not sufficient to prevent confounding when no data from the target experiment is given. The latter uses multi-view causal representation learning models to model confounding for adjustment in an observational climate application. In our paper, we also discuss conditions for identification, but we focus on a specific causal estimand, as opposed to block-identifiability of causal variables. As opposed to virtually all previous work in causal representation learning, our perspective offers clear evaluation and benchmarking potential on real-world scientific experiments – even in theoretically under-specified settings via the accuracy of the causal estimate, as also recently proposed by Yao et al. (2025).
@@ 5 Experiments
In pursuit of scientific validity, we evaluate our method (DERM) on the ISTAnt experiment (Cadei et al., 2024), and additional synthetic experiments controlling for the true causal effect.
ISTAnt We performed a similar experiment to the ISTAnt experiment on our recording platform, considering three different treatments and collecting 44 annotated 30-minute videos, and we leveraged such annotated videos to fine-tune a pretrained model generalizing the predictions to ISTAnt and then estimating the ATE by AIPW on such prediction-powered sample. We considered the AIPW estimation on the (true) ISTAnt human annotated outcomes as the ground truth ATE (approximately +40+40 short clips with $2$ clips per second, i.e., $\approx 20s$, in grooming time per video). Further details on the modeling choices, hyper-parameter, and fine-tuning are discussed in Appendix B. Figure 3 shows the differences in filming setup among the two experiments.
Figure 2: $P$-ATE inference on ISTAnt, fine-tuning different pre-trained encoders on our similar experiment, varying the training objective. Reported the 95% confidence intervals (via AIPW asymptotic normality) and compared against the same inference using the annotated outcomes. In green the correct confidence intervals (overlapping 33+%) in red the incorrect.
@@
CausalMNIST We replicate the analysis on a colored manipulation of the MNIST dataset, e.g., estimating the effect of the background color or pen color on the digit value, allowing complete control of the causal effects and uncertainty quantification by Monte Carlo simulations. We train a convolutional neural network on an experiment and we test PPCI on four similar observational studies, varying the treatment effect mechanism and experimental settings distribution, challenging the full support assumption, i.e., Equation 8. A full description of the data-generating processes and analysis is reported in Appendix C. The results are reported in Table 1, clearly showing the DERM potential to solve unsolvable instances for both ERM and invariant training (see soft shifts columns), while suffering when challenging its assumptions (see hard shifts columns).
Table 1: PPCI results on CausalMNIST varying the target experiment population (in-, out-of-distribution), the treatment effect mechanism (linear, non-linear) and challenging the full-support guarantees (soft, hard shifts). We report the ATE estimation bias and standard deviation over 50 repetitions of the data generating process.
@@ 6 Conclusion
In this paper, we formalize the problem of obtaining valid causal inference when the outcome variable is latent, and captured by complex measurements. Motivated by scientific applications, we discuss how to train an outcome predictor on similar annotated data, e.g., collected in prior experiments, and achieve valid causal conclusions on the prediction-powered target experiment. We developed the first identification and estimation results for the problem, alongside a practical constraint on the model representation to transfer causal validity under standard yet idealized assumptions. We additionally offered a novel and effective implementation of the such constraint (Deconfounded Empirical Risk Minimization), and showed that it achieves zero-shot identification of the ATE on ISTAnt by leveraging (our) similar annotated ecological experiment. Overall, our work offers a paradigm shift from the traditional Causal Representation Learning literature (Schölkopf et al., 2021), paving the way toward learning representations that power downstream causal estimates on real-world data. We think that this has the potential to play a critical role for deep learning models to accelerate the analysis of scientific data and ultimately new discoveries.
The main limitation of this work is that, in practice, it is not possible to verify whether either the conditional calibration property or the representation constraint actually holds on the target population without annotations. Beyond our key assumption that the representation transfers, we also did not discuss the model convergence on finite training datasets, so even if enforcing the representation constraint on training would transfer, we may still suffer from finite sample effects. Another limitation is neglecting several settings that are, potentially, practically relevant, e.g., not overlapping effect modifiers among similar experiments, or the violation of Equation 8. As our target experiments did not have these problems, we left them for future work. Finally, we rest on the assumption that the underlying causal inference problem is statistically identifiable, but this is necessary and can be ensured with a well-designed experimental protocol. Despite these limitations, we hope our work can inspire a new generation of practical Causal Representation Learning methodologies and create opportunities for more systematic and practically relevant benchmarking.
@@ Acknowledgments
We thank the Causal Learning and Artificial Intelligence group at ISTA for the continuous feedback on the project and valuable discussions. We thank the Social Immunity group at ISTA, particularly Jinook Oh, for the annotation program and Michaela Hoenigsberger for supporting our ecological experiment. Riccardo Cadei is supported by a Google Research Scholar Award and a Google Initiated Gift to Francesco Locatello. This research was funded in part by the Austrian Science Fund (FWF) 10.55776/COE12). It was further partially supported by the ISTA Interdisciplinary Project Committee for the collaborative project “ALED” between Francesco Locatello and Sylvia Cremer. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.